Fermi National Laboratory
 

Operation Summary for October 3, 2011 - October 5, 2011

 

Monday October 3

The day shift began with Operations monitoring the Antiproton Source (Pbar) stack (148.8E10), with the Linac, Booster, and Main Injector (MI) operating normally, and with the NuMI and MiniBooNE experiments taking beam.

During the day shift, an EE Support person found an LCW leak on a Switchyard (SY) magnet (V204) at 9:08 AM; a Fluids person began repairs. Preaccelerator personnel began work on the H- Source (the backup source) at 9:17 AM. Controls personnel began work on the TEV sector F2 vacuum crate; these areas must remain up since beam passes through sectors A and F to SY. The Fluids Group and FESS personnel began turning off the Tevatron (TEV) pond water pumps at sectors B3, C3, D3, and F3. Linac personnel began a short study period at 2:30 PM.

During the evening shift, Operations established beam to Meson FTBF experiment T-992 at 4:52 PM and disabled it at 9:12 PM. Operators reported a short downtime due to a Linac quadrupole power supply (QPS413); its panel breaker kept tripping off.

Tuesday October 4

The midnight shift began with Operations monitoring Pbar stack (142.9E10) and the protons in the Recycler (37.2E10), and with the NuMI and MiniBooNE experiments taking beam.

During the midnight shift, operators replaced a control card for QPS413. Operators reported that the power went off for MTA; the FESS Duty Electrician reset the MTA breaker. At 4:38 AM, the MiniBooNE beam permit tripped off due to bad beam positions; it reset. At 6:11 AM, the Pbar target dump cooling water inlet pressure alarmed; operators called in a Pbar Target person. At 7:24 AM, MTA lost its power again; the Duty Electrician responded. Operators reported Pbar cooling alarms due to its cryo system warming up. Operators established beam to T-992 around 7:45 AM.

During the day shift, FESS reported that the High Voltage Coordinator adjusted the MTA power breaker. At 11:11 AM, the Pbar Debuncher suffered a vacuum burst; operators closed vacuum valves and began the recovery process. Operators reported at 1:59 PM that they were having problems with SY power supply V204; operators contacted a Power Supply person. At 2:31 PM, the I- Source vacuum valve closed; it reset. A Vacuum technician reported at 2:33 PM that he was permanently closing all the vacuum valves for the A1 transfer line. The SY Power Supply person reported at 3:34 PM that V204 had a wall breaker problem, which FESS was investigating. No beam to T-992.

During the evening shift, Operations established beam to T-992 at 4:33 PM; FESS personnel had temporarily resolved the breaker problem. The MI power supply interlock loop tripped off on a MI-30 quadrupole magnet power supply ground fault indication; it reset. Operations disabled beam to T-922 at 7:03 PM.

Wednesday October 5

The midnight shift began with Operations monitoring Pbar stack (109.7) and the protons in the Recycler (17.1), and with the NuMI and MiniBooNE experiments taking beam.

During the midnight shift, operators prepared TEV enclosure B0 for an access at 12:23 AM. Operations turned off beam to NuMI at 3:30 AM for an access to move their target to a different position. Operations had beam off for MI, the Transfer Hall, and Booster by 3:57 PM; Linac personnel began beam study. A Booster RF station (BRF13) tripped off for the first of three times in 11 minutes; it reset. Linac personnel completed their study at 6:04 AM. Linac technicians began working on LRF3 at 7:02 AM and on LRF5 at 7:53 AM.

Operations Coordinator (OPSCO) Plans for Today

The plans for today has beam off in the Accelerator Complex for maintenance work and testing safety systems. The Cryo Coordinator reported that they had recovered all the helium gas and liquid to storage tanks; the nitrogen will be recovered today.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/05/2011    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory